OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -1/3 x + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (whpalmer4):

First step is to identify the slope of the line you have. It appears to be in slope-intercept form, which is: \[y=mx+b\] where \(m\) is the slope and \(b\) is the y-value of the y-intercept. The product of the slopes of two perpendicular lines will be -1. That doesn't eliminate any of your answer choices :-) You'll have to rely on your known point (-1,3) to determine which of the equations is the right answer. Only 1 will give you 3 for y if you plug in -1 for x. Which one is it?

OpenStudy (anonymous):

y = 3x - 6?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Let's check it out: y = 3(-1) + 6 y = -3 + 6 y = 3 that checks, good work!

OpenStudy (anonymous):

Thank you so much! :)